OpenStudy (anonymous):

Which equation would you get by completing the square and factoring the equation below? HELP?!?!?! x^2-16x+60=0 Answer choice: A. (x + 16)(x - 16) = 0 B. (x + 16)(x + 16) = -60 C. (x - 8)(x - 8) = 4 D. (x + 8)(x + 8) = 4

OpenStudy (anonymous):

you need answer or you need road map for solving?

OpenStudy (anonymous):

Let me see, I just need help solving this. LOL @mhmdrz91

OpenStudy (anonymous):

When completing the square, you will always take the middle term (16) (we will worry about the sign later), taking half of this (16/2 = 8), then squaring the past answer (8*8=64) now the answer you get here is what you need the final term to equal. Since the final term is 60, you must add 4 to each side of the equation: x^2 - 16x + 64 = 4 Now you take the square root of the final term of the left side (8), take the sign in front of the middle term (-), and put them in front of x, twice. (x-8)(x-8). Since we never got rid of the 4, these two will be equal to 4. { (x-8)(x+8)=4 }

OpenStudy (anonymous):

Thank's so would it be c? Unfortunately, their's no (x-8)(x+8)=4... :o

OpenStudy (anonymous):

@flixoe

OpenStudy (anonymous):

Sorry that was a typo. I did mean (x-8)(x-8)=4. C is right.

OpenStudy (anonymous):

Yay, thanks for the help!! :) @flixoe